OpenStudy (anonymous):

b. a gigabyte (gb) is 2^30 bytes. how many megabytes are in a gigabyte? how many bits are in a gigabyte?

ganeshie8 (ganeshie8):

hint : 1 megabyte = 2^20 bytes

ganeshie8 (ganeshie8):

does that help or is it confusing still hmm

OpenStudy (anonymous):

ya its

OpenStudy (anonymous):

1,024 megabytes = 1 gigabyte 210 = 1,024 bytes this is one kilobyte 220 = 1,048,576 bytes = 1 megabyte 230 = 1,073,741,824 bytes = 1 gigabyte

ganeshie8 (ganeshie8):

glad to hear :)

ganeshie8 (ganeshie8):

omg ! you got them everything in numbers

ganeshie8 (ganeshie8):

1024 = 2^10

ganeshie8 (ganeshie8):

let me show u how to get the answer in powers of 2

ganeshie8 (ganeshie8):

1 gigabyte = \(2^{30}\) bytes = \(2^{10+20}\) bytes = \(2^{10} . 2^{20} \) bytes = \(2^{10} \) megabytes

OpenStudy (anonymous):

oh lol

ganeshie8 (ganeshie8):

hmm